OpenStudy (anonymous):

let f(x) = 2x^2-5x+9 a.Find slope of tangent line at x=2 b. Find an equation of the tangent line at x=2

OpenStudy (anonymous):

I'm assuming you're in calculus? If so, the slope is the first derivate of the function.

OpenStudy (anonymous):

Yes calc

Parth (parthkohli):

The slope of your tangent line is the derivative.

OpenStudy (anonymous):

I have no idea how to set this up

OpenStudy (anonymous):

I have done problems like this though so I "know" how to do them and Ive done them before! thats whats frustrating lol

OpenStudy (anonymous):

you can find the equation of the tangent line by doing y = mx + b form where y is the function of your tangent line, m = f'(2) the derivative at x=2, x is your independent varaible, and b is the y-int of the tangent line.

Parth (parthkohli):

I'm telling you how to do (a). 1: Find \(\rm f'(x)\). Here, it is \(\rm f'(x) = 4x - 5\). 2: Evaluate \(\rm f'(x)\) for \(\rm x = 2\). Here, it is \(\rm 4(2) - 5 = 3\)

OpenStudy (anonymous):

a. Calculate the derivate and calculate the value for x = 2 f'(x) = 4x -5 f'(2) = 4x2 -5 = 3 b. Calculate a point using f(x) and use that with the slope to calculate the formula for the line f(2) = 7 point (2,7) slope 3 y = ax + b 7 = 3 times 2 + b 7 = 6 + b b = 1 y = 3x + 1

OpenStudy (anonymous):

Ok thanks everyone Im going through this now

OpenStudy (anonymous):

where did the 4x-5 come from ? for the derivative?

OpenStudy (anonymous):

it's the derivate of your function -> 2x^2 => 4x -5x => -5

OpenStudy (anonymous):

I guess I don't know how to do that

OpenStudy (anonymous):

You had best google it or open your book

OpenStudy (anonymous):

already there, problem is Im sick and this cold medicine is blocking my memory right now I can't form a thought properly thats why I can't do this

OpenStudy (anonymous):

Well, i can't really help you understand either then, use the answer i gave you, get better and study :)

OpenStudy (anonymous):

ok the power rule, jesus :P I got it

OpenStudy (anonymous):

ghehe

OpenStudy (anonymous):

thanks everyone